OpenStudy (anonymous):

How many solutions does the equation 5x + 10 = 5x − 8 have? Zero One Two Infinitely many

OpenStudy (anonymous):

i give out medals for the best or the right answers

OpenStudy (anonymous):

To find the answer to this question you have to combine like terms. To do this you are going to take the 5x from the right side of the equation and subtract it from the 5x on the left side (you subtract because you also have to take the 5x off the right side of the equation and to do this you have to subtract to cancel out). After you subtract the 5x you see that your equation is 10=-8 you can no longer do anything with this equation and you do not see what x equals, this means the answer is zero.

OpenStudy (anonymous):

okay hold on a second and thank you for answering

OpenStudy (anonymous):

can you help me with two more please

OpenStudy (anonymous):

sure :D

OpenStudy (anonymous):

What is the value of x in the equation 3(2x + 4) = −6? −3 1 12 19

OpenStudy (anonymous):

To find this answer all you have to do is distribute the 3 into the equation so you do 3*2x=6x and then you do 3*4=12 so your new equation is 6x+12=-6. Since you want to find out what x equals you have to get it alone on one side of the equation. To do this you subtract 12 from both sides of the equation to get the new equation of 6x=-18 now you divide 6 from both sides of the equation to get x=-3